University of Minnesota Lab NFC

Badger LMS

Problems and Shutdowns

Problem  18 Dec K3 Sputterer AJA1 waiting for power supply parts -> Power supply is at vendor and we are still waiting for parts before it can be repaired and shipped back. will update when we have more information. olms0025
Problem  8 May P5 Ebeam Litho EBPG5000+ Vistec use only 100 nA or lower beams -> column is out of alignment, so use only 100 nA or lower beams. rober074
Problem  16 May PAN 185 Stabino Zeta Potential Analyzer Stabino II Requires Calibration and Diagnostics Testing -> All beaker and piston combinations require calibration and follow-up diagnostic testing to confirm accuracy and precision in its analysis of Zeta Potential for colloidal suspensions. Beakers are labeled 1, 2, Micro. Pistons are labeled 200 um, 400 um, and 1 mL Calibrations required for each beaker and piston combo include: Single Point Zeta Potential (Ludox TM-40 silica, -38 mV) Multi-Point Zeta Potential (Ludox TM-40 silica, -38 mV) Acid Base Titrations - pH meter (Triple point using buffers 4.0, 7.0, and 10.0) Streaming Potential (P-DADMAC at 0.0025N) Please note that this is a shared tool workstation with the Microtrac DLS Analyzer. Work should be scheduled around DLS tool appointments. I will be able to perform these calibrations and diagnostic testing on Monday, May 19th at 9 am. Please allow up to 2-3 hours for completion. Thanks, - Matthew Johnson NanoBio Lab Support Specialist Minnesota Nano Center - PAN-140 115 Union Street SE Minneapolis, MN 55455 joh14999 at umn.edu C: (612)306-2796 joh14999
Problem  19 May K1 AFM DI3000 needs replacement video monitor -> need vga monitor and cable if available for a successful start up broug001
Problem  21 May K3 Etcher XeF2 Xactix tool is available -> Tool is available for use but is shutdown to preserve pump life. If you need this tool, just email MNC staff and we'll have it up in 48 hours. olms0025
Problem  22 May P3 Wet Bench HF Process Beaker of BOE Sitting Out -> There has been a beaker of BOE sitting at this bench for many months that needs to be disposed of, as this feels rather unsafe. It seems that over the months it has also been filled up more so it is now close to full. nilso037
Problem  23 May P3 Sputterer AJA2 G2 plasma doesn't turn on kim02384
Problem  23 May P3 Sputterer AJA2 Gun 2 seems to be shorted -> I took gold out of gun, looked for flakes, reassembled, but still seems shorted. Won't ight until problem is found. rober074
Problem  6 Dec K1 Oxidation Tube 14 Gate Oxide Tylan tycom central control console reliability pm -> doing biannual pm on Tycom CCC by disassembling and checking power supply voltages and cleaning PC board edge connections with contact cleaner as a reliability pm broug001
Problem  28 Nov KA2 CVIV testing IV - Computer issues -> A replacement computer is being worked on. The CV and IV HP systems can be used in manual mode. Without computer no data files can be saved. whipp003
Problem  14 Mar K1 ALD Savannah Ultratech Current Films update -> Precursors were rearragnged Al2O3, HfO2 and SnO are now available. Do not use SiO2. jorex023
Shutdown 23 Apr P1 Ellipsometer Rudolph shutdown update -> "System was making random crackling sounds and found the laser was not on. Turned system power off using key on the side." This is still in queue, will probably not get to this for several weeks. If you need the tool urgently, contact Brian Olmsted olms0025
Shutdown 16 Aug KA1 Si Nanowire Deposition exhaust off cibuzar
Shutdown 4 Feb K1 Oxidation Tube 15 General Tylan Process person wants to do some content analysis test runs in the near future -> due to T15 having been used for special project use since 2020. broug001
Shutdown 3 Oct K1 Oxidation Tube 14 Gate Oxide Tylan serioal io error and ot TC error and torch controller needs to heat -> no steam oxidations until torch temp exceeds 650C to satisfy H2 temp interlock broug001
Shutdown 2 Apr P3 HD PCVD Plasmatherm chamber components are cleaned. needs re-assembly. -> We have a pump ready to go. Need to Tee this into exhaust facilities and then re-assemble chamber components. Will start working on this again soon, as other tasks are nearing completion. olms0025
Shutdown 30 Apr K1 Wafer Bonder SB6 Karl Suss had tool not moved error and will install the new air cylinders -> when time permits broug001
Shutdown 14 May K1 RTA Solaris 150 sent back to mfgr -> The RTA was sent back to the manufacturer to repair a persistent problem. No return ETA yet, but they will start working on this tomorrow. olms0025
Shutdown 21 May K1 Ellipsometer Spectroscopic Woolam new power supply arrived -> New power supply has arrived. needs to be installed. olms0025
Shutdown 21 May K3 Evaporator Temescal base pressure was 3.4e-7 except at pocket 3 it hopped to 1e-4T -> RM and TB here assessing ebgun hearth alignment and replacing seals and precision aligning hearth before reassembly and loading a tooling factor Aluminum deposition test; We will check 0-10V output calibration 0 to 100 percent on dep controller before depositing broug001
Shutdown 23 May P1 UHV Deposition AJA chamber pressure near -8T range needing bakeout -> Lamps are extended, both heater pads are installed. Chamber needs bakeout. single head qcm head is installed on qcm channel 1 broug001

Equipment In Use

P3 Evaporator CHA ren00084

Recent Cancellations

kim02384 23 May 14:00 15:00 P3 Sputterer Gold Target AJA2
kim02384 23 May 14:00 15:00 P3 Sputterer AJA2
espe0124 23 May 14:00 17:00 P5 Ebeam Litho EBPG5000+ Vistec

Today's Reservations

ren00084 10:00 13:15 P5 Ebeam Litho EBPG5000+ Vistec
ren00084 13:15 14:15 P5 Ebeam Litho EBPG5000+ Vistec
ren00084 14:15 15:15 P5 Ebeam Litho EBPG5000+ Vistec

Reservations for Tomorrow

hoque028 11:00 12:30 P5 Ebeam Litho EBPG5000+ Vistec
hoque028 13:00 15:30 P3 Evaporator CHA
hoque028 17:30 19:00 P5 Ebeam Litho EBPG5000+ Vistec
hoque028 19:00 21:30 P3 Evaporator CHA

Comments

P5 Ebeam Litho EBPG5000+ Vistec 28 Apr beam files redefined 4-25-2025 -> beam files from 120 pA to 200 nA redfined 4-25-2025. rober074
K1 Ellipsometer LSE Gaertner 30 Apr obtaining specific wafers -> to perform calibration procedure in user manual broug001
KA1 Dicing Saw DAC 552 6 May silicon blade installed -> the silicon blade is currently installed. rober074
K3 Evaporator Temescal 21 May beadblasting temescal 6 pocket hearth -> to start off with a clean bare hearth broug001
K3 Sputterer AJA1 23 May Gun 2 not working bc no power supply -> Gun 2 not working bc no power supply. Supply is out for repair. rober074
P3 Sputterer AJA2 23 May all guns returned to orig. targets -> All guns have all been returned to their original targets. Al, Au, SiO2, Cu, Ti, WTi. rober074
P3 Evaporator CHA 24 May Noticed IG2 setpoint had been changed -> Noticed IG2 setpoint for pumpdown had been changed to 4e-6 from 1.3e-6 noted in SOP. Unsure if this was intentional or not rigge016
P3 Evaporator CHA 24 May Ti source is possibly missing -> Solid Ti source is possibly missing. An unidentified source was left in pocket 4 by a previous user but I cannot verify from run log sheet if this is Ti or not. Unidentified source is on a napkin by the scale. rigge016